# Identifying Compton-thick AGNs in the COSMOS

### v1

@@ Abstract
Context. Compton-thick active galactic nuclei (CT-AGNs), defined by column density , are so heavily absorbed that their X-ray emission is often feeble, even undetectable by X-ray instruments. Nevertheless, their radiation is expected to be a substantial contributor to the cosmic X-ray background (CXB), predicting that CT-AGNs should comprise at least 30% of the total AGN population.
Aims. In the Cosmological Evolution Survey (COSMOS), the identified CT-AGN fraction falls far below theoretical expectations, indicating that a substantial population of CT-AGNs is hidden due to their low photon counts or their flux below the current flux limits of X-ray instruments. This work focuses on identifying CT-AGNs hidden in mid-infrared (MIR)-selected AGNs.
Methods. First, we selected a sample of 1,104 MIR-selected AGNs that were covered but individually undetected by X-ray. Next, we reduced the X-ray data in the COSMOS and analyzed multiwavelength data in our sample to derive the key physical parameters required for CT-AGN identification.
Results. Using MIR diagnostics, we first find out 7 to 23 CT-AGN candidates. Their subsequent X-ray stacking analysis reveals a clear detection at significance in the soft band and only significance in the hard band. We fit the stacked soft- and hard-band fluxes with a physical model and confirm that these sources are absorbed by Compton-thick material. However, CT-AGNs constitute only 2.1% (23/1104) of our sample, significantly below the fraction predicted by CXB synthesis models, indicating that a considerable population of CT-AGNs remains missed by our selection. A comparison of host-galaxy properties between CT-AGNs and non-CT-AGNs reveals no significant differences.
@@
Active galactic nuclei (AGNs) are defined as galaxies hosting supermassive black holes (SMBHs, ) that have an Eddington ratio111The Eddington ratio is defined as the ratio of the bolometric luminosity of the AGN to the Eddington luminosity, which represents the balance between the gravitational force and the radiation pressure. For a supermassive black hole with mass , the Eddington luminosity can be approximated as: exceeding the limit of (Netzer, 2015). These systems are powered by the accretion of surrounding matter through the SMBHs, thereby releasing their gravitational potential energy. Their emissions cover the entire electromagnetic spectrum, exhibiting distinct radiation characteristics at varying distances from the central SMBHs. The AGN unification model suggests that the radiation across different bands originates from their specific substructures (Antonucci, 1993; Urry & Padovani, 1995; Netzer, 2015). For instance, the accretion disk near the SMBH is heated by the viscous dissipation of gravitational potential energy during mass accretion, producing a broad ultraviolet (UV)-to-optical continuum spectrum (e.g., Richstone & Schmidt, 1980). The hot corona, hovering above the accretion disk, generates luminous X-ray emission through inverse Compton up-scattering of UV–optical photons originating from the accretion disk (Liang, 1979; Haardt & Maraschi, 1991). The dusty torus, situated outside the corona and accretion disk, absorbs the intense radiation from the central engine and reprocesses it into thermal emission at longer mid-infrared (MIR) band (e.g., Antonucci, 1993). Owing to the evolutionary stages of AGNs and the viewing angle effects, the majority of AGNs are obscured by large amounts of dust and gas (e.g., Hickox & Alexander, 2018). The high penetration ability of X-ray radiation, combined with the bright X-ray emissions from AGNs, makes X-rays an essential tool for distinguishing AGNs from other extragalactic sources and for investigating the obscured material within AGNs (e.g., Xue, 2017). The neutral hydrogen column density () is a key parameter for describing the obscured AGNs in the X-ray band. The obscured AGNs exhibiting a column density of are classified as Compton-thick AGNs (CT-AGNs; e.g., Comastri, 2004).
@@
The cosmic X-ray background (CXB) is the radiation from the superposition of all discrete X-ray point sources, of which AGNs account for the vast majority (Mushotzky et al., 2000; Moretti et al., 2003; Luo et al., 2017). The most prominent observational feature of the CXB spectrum is a pronounced hump, marking the peak of its energy density, around 30 keV (Revnivtsev et al., 2003; Churazov et al., 2007; Frontera et al., 2007; Ajello et al., 2008). Over the past several decades, theoretical modeling of the CXB spectrum has repeatedly demonstrated that only AGN population synthesis models incorporating a substantial fraction of CT-AGNs can successfully reproduce the complete spectral shape of the CXB (e.g., Gilli et al., 2007; Treister et al., 2009; Burlon et al., 2011; Akylas et al., 2012; Ueda et al., 2014; Ananna et al., 2019; Georgantopoulos et al., 2025). To date, the fraction of CT-AGNs remains uncertain and is a subject of hot debate. At low redshifts, such as in the local Universe, AGN population synthesis models predict that CT-AGNs account for 20–30% of the total AGN population detected by hard X-ray surveys such as Swift-BAT (Burlon et al., 2011; Ueda et al., 2014; Ricci et al., 2015). The denser atomic and molecular gas content in high-redshift galaxies (e.g., Carilli & Walter, 2013), relative to the local Universe, increases the likelihood that AGNs hosted in these environments are heavily obscured, thereby leading to a higher expected fraction of CT-AGNs. For instance, Buchner et al. (2015) estimated that CT-AGNs account for of the total AGN population in X-ray deep-field surveys, while Ananna et al. (2019) found that this fraction increases to for AGNs within the redshifts .
@@
Nevertheless, attaining the theoretically anticipated fraction of CT-AGNs within X-ray surveys remains a significant challenge. The X-ray radiation of CT-AGNs is typically extremely faint, so their photon counts are low, even undetectable. These characteristics make CT-AGNs the most difficult population to detect in X-ray surveys (e.g., Matt et al., 2000; Comastri et al., 2015; Padovani et al., 2017). In particular, for sources with column densities , even in the local Universe, X-ray flux-limited surveys are biased against detecting CT-AGNs (e.g., Ricci et al., 2015; Koss et al., 2016; Annuar et al., 2025; Boorman et al., 2025). X-ray surveys typically report an observed fraction of CT-AGNs below 15% (e.g., Ricci et al., 2017b; Masini et al., 2018; Georgantopoulos & Akylas, 2019), which is significantly lower than theoretical model predictions, meaning that a large number of CT-AGNs remain unidentified or undiscovered. These missing CT-AGNs may still be hidden in AGNs that exhibit low photon counts (e.g., Lambrides et al., 2020) or that have not been detected by X-ray surveys (e.g., Carroll et al., 2021). To identify the missing CT-AGNs in the local Universe, the X-ray astronomy group at Clemson University and the extragalactic astronomy group at the INAF-OAS Bologna and INAF-IASF Palermo initiated a targeted project in 2017 to actively search for CT-AGNs, known as the Clemson-INAF CT-AGN Project (Marchesi et al., 2018). The project used a volume-limited sample of CT-AGNs with NuSTAR spectra, achieving a CT-AGN fraction of within the redshift (Torres-Albà et al., 2021). Similarly, a series of studies have focused on searching for the missing CT-AGNs in X-ray deep-field surveys. For example, Guo et al. (2021) identified eight missing CT-AGNs by using the multiwavelength techniques in the Chandra Deep Field-South (CDFS) survey. Zhang et al. (2025) used machine learning algorithms to identify 67 missing CT-AGNs, thereby raising the fraction of CT-AGNs to 20% in the CDFS survey. Numerous efforts have been made to search for the missing CT-AGNs in X-ray surveys, yet the theoretical expectations remain unmet.
@@
The Cosmic Evolution Survey (COSMOS; e.g., Scoville et al., 2007) is a deep, multiwavelength survey covering a relatively large area of 2 deg2, aimed at measuring the evolution of galaxies and AGNs. The X-ray observations of this field are also crucial for resolving the CXB (e.g., Buchner et al., 2015). The Chandra X-ray telescope has covered the entire COSMOS, achieving a relatively deep average exposure (e.g., Civano et al., 2016; Marchesi et al., 2016). However, the observed fraction of CT-AGNs in the COSMOS is significantly lower than theoretical expectations. For example, Lanzuisi et al. (2018) successfully identified 67 CT-AGNs from a sample of 1855 AGNs with high (¿30) photon counts through X-ray spectral fitting. To find out the hidden CT-AGNs in the COSMOS, it is essential to establish and execute a series of specific tasks. Guo et al. (2025) have completed the first task and successfully identified 18 CT-AGNs. In this work, we aim to execute and complete the second task, focusing on identifying CT-AGNs among MIR-selected AGNs. This paper is structured as follows. Sect. 2 presents the data used to select AGNs via MIR colors and details the construction of our final sample. In Sect. 3, we describe the Chandra observations covering COSMOS and their data reduction. Sect. 4 presents the multiwavelength photometric data for the sample and the corresponding analysis. Sect. 5 presents our CT-AGN diagnostics, briefly discusses the missing CT-AGNs in our sample and the properties of CT-AGN host galaxies . A brief summary is presented in Sect. 6. We adopt a concordance flat -cosmology with , , and (Planck Collaboration et al., 2020). All quoted uncertainties correspond to the 1 (68%) confidence level unless stated otherwise, while upper limits are given at the 3 level.
@@ 2.1 MIR data Collation
To construct a high-quality MIR-selected AGN sample, we first systematically collated the MIR data. The primary data source was the COSMOS2020 FARMER catalog (Weaver et al., 2022), which provides Spitzer/IRAC four-band (3.6 m, 4.5 m, 5.8 m, 8.0 m) photometric data. To meet the requirements for MIR AGN selection, the following constraints were applied to the MIR photometric data:
(1).
Sources with redshifts between 0.1 and 2 were selected.
(2).
Sources with IRAC magnitudes not less than the detection limit were chosen to ensure data reliability.
(3).
The flux-to-error ratio in each band was required to be greater than 5 to ensure high signal-to-noise ratios.
In the COSMOS2020 FARMER catalog, a total of 56,831 sources satisfy the above constraints. We have collected the four band flux densities of Spitzer/IRAC for use in the next section.
@@ 2.2 Selection of AGN sample
The unique colors of AGNs in the MIR band, especially between 3–5 m, make their spectra appear redder than those of normal galaxies. Therefore, many studies use this feature to distinguish AGN from normal galaxies and identify them (e.g., Lacy et al., 2004, 2007, 2013; Donley et al., 2012; Assef et al., 2013; Chang et al., 2017). For example, the IRAC color-color diagrams most commonly used for AGN selection were defined by Lacy et al. (2004, 2007, 2013) and Donley et al. (2012). In this work, we adopted the AGN selection criterion of Donley et al. (2012), which is more stringent than others and yields a highly pure and reliable AGN sample. Fig. 1 presents an IRAC color-color diagram, plotting against . The sources within the red dashed lines are AGNs selected using the criteria from Donley et al. (2012), totaling 1,545. Among them 252 are X-ray sources, of which nine have been classified as CT-AGNs by Lanzuisi et al. (2018) and Guo et al. (2025).
@@
(1).
Remove known X-ray sources from the MIR-selected AGNs.
(2).
Eliminate sources lacking X-ray observations (i.e., those with exposure time = 0).
(3).
@@
After this screening, our final AGN sample comprises 1,104 sources. Cols. 1–3 of Table 2 list their IDs, RA, and DEC. The photometric redshifts of our AGN sample are taken from the COSMOS2020 FARMER catalog, measured by LePhare using the galaxy templates. For the bright sources, Weaver et al. (2022) find a normalized median absolute deviation of only 0.009 in their photometric redshifts; even in the faintest bin (25 mag ¡ i ¡ 27 mag) the normalized median absolute deviation is 0.036. Col. 4 of Table 2 lists their photometric redshifts.
@@ 3.1 Overview of Chandra’s Observations in COSMOS
To identify (or probe) AGNs with varying redshifts in the COSMOS and investigate black hole accretion history as well as the co-evolution of AGNs with their host galaxies, Elvis et al. (2009) conducted Chandra observations during 2006–2007, accumulating a total exposure of 1.8 Ms. These observations covered the central 0.5 deg2 with an effective exposure of approximately 160 ks and an outer 0.4 deg2 area with an effective exposure of approximately 80 ks. Subsequently, Civano et al. (2016) added further Chandra observations during 2013–2014, increasing the total exposure to 2.8 Ms. These additional observations covered the remaining area (the outer 0.4 deg2 from Elvis et al., 2009) with an effective exposure of approximately 80 ks and increased the effective exposure over the central 1.5 deg2 to approximately 160 ks. As a result, the combined observations totaling approximately 4.6 Ms covered a region of 2.2 deg2 in the COSMOS. For detailed information on observation files, including observation and exposure times.
@@ 3.3 X-ray luminosity upper limit estimations
Our MIR-selected AGNs are individually undetected in X-rays, their X-ray luminosities cannot be measured directly. Consequently, we quote only upper limits, calculated with the equation (Alexander et al., 2003),
@@
where is the upper limit of the count rate at the undetected source and CF is the energy conversion factor, set at 3.06 (Civano et al., 2016). The upper limit of the count rate for undetected sources is calculated by
where represents the effective exposure time at the source location , denotes the upper limit on the background counts measured within a -radius aperture centered on the source. The upper limit on the background counts is calculated using Equation 9 of Gehrels (1986), giving
where represents the 0.5–8 keV background counts measured within a -radius aperture centecalibrered on the source, after correction for the energy-encircled fraction. The background counts for AGNs are determined by summing all values from a radius circle centered at their respective coordinates in the modeled background map. Similarly, the exposure times for the AGNs are determined by reading the values from the exposure map at their respective coordinates. Fig. 2 shows the coordinates of MIR-selected AGNs on the exposure map and their corresponding exposure times. Cols. 5–6 of Table 2 give the aperture background counts and exposure time for each source.
Applying the calculation process described above, we derived X-ray luminosity upper limits for the entire MIR-selected AGN sample. Cols. 7–8 of Table 2 list the upper limits of flux density and luminosity in the 2–10 keV band.

### v2

# Identifying Compton-thick active galactic nuclei in the COSMOS

@@ Abstract
Context. Compton-thick active galactic nuclei (CT-AGNs), defined by a column density of , are so heavily absorbed that their X-ray emission is often feeble and can even be undetectable by X-ray instruments in some cases. Nevertheless, their radiation is expected to be a substantial contributor to the cosmic X-ray background (CXB), predicting that CT-AGNs would comprise at least 30% of the total AGN population.
Aims. Cosmological Evolution Survey (COSMOS) reported that the identified CT-AGN fraction falls far below theoretical expectations, indicating that a substantial population of CT-AGNs is hidden due to their low photon counts or due to their flux lying below the current flux limits of X-ray instruments. This work focuses on identifying CT-AGNs hidden among mid-infrared (MIR) selected AGNs.
Methods. First, we selected a sample of 1,104 MIR-selected AGNs that were covered, but individually undetected by X-ray. Next, we reduced the X-ray data in the COSMOS and analyzed multiwavelength data in our sample to derive the key physical parameters required for the CT-AGN identification.
Results. Using MIR diagnostics, we found 7 to 23 CT-AGN candidates. Their subsequent X-ray stacking analysis revealed a clear detection at significance in the soft band and only a significance in the hard band. We fit the stacked soft- and hard-band fluxes with a physical model and confirm that these sources are absorbed by Compton-thick material. However, CT-AGNs ultimately constituted only 2.1% (23/1104) of our sample, significantly below the fraction predicted by CXB synthesis models. This indicates that a considerable population of CT-AGNs remains missed by our selection. A comparison of host-galaxy properties between CT-AGNs and non-CT-AGNs reveals no significant differences.
@@
Active galactic nuclei (AGNs) are defined as galaxies hosting supermassive black holes (SMBHs, ) that have an Eddington ratio111The Eddington ratio is defined as the ratio of the bolometric luminosity of the AGN to the Eddington luminosity, which represents the balance between the gravitational force and the radiation pressure. For a supermassive black hole with mass , the Eddington luminosity can be approximated as exceeding the limit of (Netzer, 2015). These systems are powered by the accretion of surrounding matter through the SMBHs, thereby releasing their gravitational potential energy. Their emissions cover the entire electromagnetic spectrum, exhibiting distinct radiation characteristics at varying distances from the central SMBHs. The AGN unification model suggests that the radiation across different bands originates from their specific substructures (Antonucci, 1993; Urry & Padovani, 1995; Netzer, 2015). For instance, the accretion disk near the SMBH is heated by the viscous dissipation of gravitational potential energy during mass accretion, producing a broad ultraviolet (UV) to optical continuum spectrum (e.g., Richstone & Schmidt, 1980). The hot corona, hovering above the accretion disk, generates luminous X-ray emission through inverse Compton up-scattering of UV–optical photons originating from the accretion disk (Liang, 1979; Haardt & Maraschi, 1991). The dusty torus, situated outside the corona and accretion disk, absorbs the intense radiation from the central engine and reprocesses it into thermal emission at longer mid-infrared (MIR) band (e.g., Antonucci, 1993). Owing to the evolutionary stages of AGNs and the viewing angle effects, the majority of AGNs are obscured by large amounts of dust and gas (e.g., Hickox & Alexander, 2018). The high penetration ability of X-ray radiation, combined with the bright X-ray emissions from AGNs, makes X-rays an essential tool for distinguishing AGNs from other extragalactic sources and for investigating the obscured material within AGNs (e.g., Xue, 2017). The neutral hydrogen column density () is a key parameter for describing the obscured AGNs in the X-ray band. The obscured AGNs exhibiting a column density of are classified as Compton-thick AGNs (CT-AGNs; e.g., Comastri, 2004).
@@
The cosmic X-ray background (CXB) is the radiation from the superposition of all discrete X-ray point sources and AGNs account for the vast majority of these sources (Mushotzky et al., 2000; Moretti et al., 2003; Luo et al., 2017). The most prominent observational feature of the CXB spectrum is a pronounced hump, marking the peak of its energy density, around 30 keV (Revnivtsev et al., 2003; Churazov et al., 2007; Frontera et al., 2007; Ajello et al., 2008). Over the past several decades, theoretical modeling of the CXB spectrum has repeatedly demonstrated that only AGN population synthesis models incorporating a substantial fraction of CT-AGNs can successfully reproduce the complete spectral shape of the CXB (e.g., Gilli et al., 2007; Treister et al., 2009; Burlon et al., 2011; Akylas et al., 2012; Ueda et al., 2014; Ananna et al., 2019; Georgantopoulos et al., 2025). To date, the fraction of CT-AGNs remains uncertain and is a subject of hot debate. At low redshifts, as in the case of the Local Universe, AGN population synthesis models predict that CT-AGNs account for 20–30% of the total AGN population detected by hard X-ray surveys such as Swift-BAT (Burlon et al., 2011; Ueda et al., 2014; Ricci et al., 2015). The denser atomic and molecular gas content in high-redshift galaxies (e.g., Carilli & Walter, 2013), relative to the Local Universe, increases the likelihood that AGNs hosted in these environments are heavily obscured, thereby leading to a higher expected fraction of CT-AGNs. For instance, Buchner et al. (2015) estimated that CT-AGNs account for of the total AGN population in X-ray deep-field surveys, while Ananna et al. (2019) found that this fraction increases to for AGNs throughout redshift .
@@
Nevertheless, attaining the theoretically anticipated fraction of CT-AGNs within X-ray surveys remains a significant challenge. The X-ray radiation of CT-AGNs is typically extremely faint, so their photon counts are low and often even undetectable. These characteristics make CT-AGNs the most difficult population to detect in X-ray surveys (e.g., Matt et al., 2000; Comastri et al., 2015; Padovani et al., 2017). In particular, for sources with column densities , even in the Local Universe, X-ray flux-limited surveys are biased against detecting CT-AGNs (e.g., Ricci et al., 2015; Koss et al., 2016; Annuar et al., 2025; Boorman et al., 2025). X-ray surveys typically report an observed fraction of CT-AGNs below 15% (e.g., Ricci et al., 2017b; Masini et al., 2018; Georgantopoulos & Akylas, 2019), which is significantly lower than theoretical model predictions, meaning that a large number of CT-AGNs remain unidentified or undiscovered. These missing CT-AGNs may still be hidden in AGNs that exhibit low photon counts (e.g., Lambrides et al., 2020) or that have not been detected by X-ray surveys (e.g., Carroll et al., 2021). To identify the missing CT-AGNs in the Local Universe, the X-ray astronomy group at Clemson University and the extragalactic astronomy group at the INAF-OAS Bologna and INAF-IASF Palermo initiated a targeted project in 2017 to actively search for CT-AGNs, known as the Clemson-INAF CT-AGN Project (Marchesi et al., 2018). The project used a volume-limited sample of CT-AGNs with NuSTAR spectra, achieving a CT-AGN fraction of within the redshift of (Torres-Albà et al., 2021). Similarly, a series of studies have focused on searching for the missing CT-AGNs in X-ray deep-field surveys. For example, Guo et al. (2021) identified eight missing CT-AGNs by using the multiwavelength techniques in the Chandra Deep Field-South (CDFS) survey. Zhang et al. (2025) used machine learning algorithms to identify 67 missing CT-AGNs, thereby raising the fraction of CT-AGNs to 20% in the CDFS survey. Numerous efforts have been made to search for the missing CT-AGNs in X-ray surveys, yet the theoretical expectations remain unmet.
@@
The Cosmic Evolution Survey (COSMOS; e.g., Scoville et al., 2007) is a deep, multiwavelength survey covering a relatively large area of 2 deg2, aimed at measuring the evolution of galaxies and AGNs. The X-ray observations of this field are also crucial for resolving the CXB (e.g., Buchner et al., 2015). The Chandra X-ray telescope has covered the entire COSMOS, achieving a relatively deep average exposure (e.g., Civano et al., 2016; Marchesi et al., 2016). However, the observed fraction of CT-AGNs in the COSMOS is significantly lower than theoretical expectations. For example, Lanzuisi et al. (2018) successfully identified 67 CT-AGNs from a sample of 1855 AGNs with high (¿30) photon counts through X-ray spectral fitting. To find out the hidden CT-AGNs in the COSMOS, it is essential to establish and execute a series of specific tasks. Guo et al. (2025) completed the first task and successfully identified 18 CT-AGNs. In this work, we aim to execute and complete the second task, focusing on identifying CT-AGNs among MIR-selected AGNs. This paper is structured as follows. Section 2 presents the data used to select AGNs via MIR colors and details the construction of our final sample. In Sect. 3 we describe the Chandra observations covering COSMOS and their data reduction. Section 4 presents the multiwavelength photometric data for the sample and the corresponding analysis. Section 5 presents our CT-AGN diagnostics, along with a brief discussion of the missing CT-AGNs in our sample and the properties of CT-AGN host galaxies. A brief summary is presented in Sect. 6. We adopted a concordance flat -cosmology with , , and (Planck Collaboration et al., 2020). All quoted uncertainties correspond to the 1 (68%) confidence level unless stated otherwise, while upper limits are given at the 3 level.
@@ 2.1 MIR data collation
To construct a high-quality MIR-selected AGN sample, we systematically collated the MIR data. The primary data source was the COSMOS2020 FARMER catalog (Weaver et al., 2022), which provides Spitzer/IRAC four-band (3.6 m, 4.5 m, 5.8 m, and 8.0 m) photometric data. To meet the requirements for the sample selection, the following constraints were applied to the MIR photometric data:
Sources with redshifts between 0.1 and 2 were selected;
Sources with IRAC magnitudes over the detection limit were chosen to ensure data reliability;
The flux-to-error ratio in each band was required to be greater than 5 to ensure a high signal-to-noise ratios.
In the COSMOS2020 FARMER catalog, a total of 56,831 sources satisfy the above constraints. We collected the four band flux densities of Spitzer/IRAC for the steps described in the next section.
@@ 2.2 Selection of AGN sample
The unique colors of AGNs in the MIR band, especially between 3–5 m, make their spectra appear redder than those of normal galaxies. Therefore, many studies use this feature to distinguish AGNs from normal galaxies and identify them (e.g., Lacy et al., 2004, 2007, 2013; Donley et al., 2012; Assef et al., 2013; Chang et al., 2017). For example, the IRAC color-color diagrams most commonly used for AGN selection were defined by Lacy et al. (2004, 2007, 2013) and Donley et al. (2012). In this work, we adopted the AGN selection criterion of Donley et al. (2012), which is more stringent than others and yields a highly pure and reliable AGN sample. Figure 1 presents an IRAC color-color diagram, plotting against . The sources within the red dashed lines are AGNs selected using the criteria from Donley et al. (2012), totaling 1,545. Among them 252 are X-ray sources, of which nine have been classified as CT-AGNs by Lanzuisi et al. (2018) and Guo et al. (2025).
@@
Remove known X-ray sources from the MIR-selected AGNs;
Eliminate sources lacking X-ray observations (i.e., those with exposure time = 0);
@@
After this screening, our final AGN sample comprises 1,104 sources. Columns 1–3 of Table 2 list their IDs, RA, and DEC. The photometric redshifts of our AGN sample are taken from the COSMOS2020 FARMER catalog, measured by LePhare using the galaxy templates. For the bright sources, Weaver et al. (2022) find a normalized median absolute deviation of only 0.009 in their photometric redshifts; even in the faintest bin (25 mag ¡ i ¡ 27 mag) the normalized median absolute deviation is 0.036. Column 4 of Table 2 lists their photometric redshifts.
@@ 3.1 Overview of Chandra’s observations in COSMOS
To identify (or probe) AGNs with varying redshifts in the COSMOS and investigate black hole accretion history, as well as the co-evolution of AGNs with their host galaxies, Elvis et al. (2009) conducted Chandra observations during 2006–2007, accumulating a total exposure of 1.8 Ms. These observations covered the central 0.5 deg2 with an effective exposure of approximately 160 ks and an outer 0.4 deg2 area with an effective exposure of approximately 80 ks. Subsequently, Civano et al. (2016) added further Chandra observations during 2013–2014, increasing the total exposure to 2.8 Ms. These additional observations covered the remaining area (the outer 0.4 deg2 from Elvis et al., 2009) with an effective exposure of approximately 80 ks and increased the effective exposure over the central 1.5 deg2 to approximately 160 ks. As a result, the combined observations totaling approximately 4.6 Ms covered a region of 2.2 deg2 in the COSMOS. For more detailed information on observation files, including observation and exposure times, we refer to Elvis et al. (2009) and Civano et al. (2016).
@@ 3.3 X-ray luminosity upper limit estimations
Our MIR-selected AGNs are individually undetected in X-rays, their X-ray luminosities cannot be measured directly. Consequently, we quote only upper limits, calculated with the equation (Alexander et al., 2003):
@@
where is the upper limit of the count rate at the undetected source and CF is the energy conversion factor, set to 3.06 (Civano et al., 2016). The upper limit of the count rate for undetected sources is calculated by
where represents the effective exposure time at the source location , denotes the upper limit on the background counts measured within a radius aperture centered on the source. The upper limit on the background counts can be calculated using Eq. 9 of Gehrels (1986), giving
where represents the 0.5–8 keV background counts measured within a -radius aperture centecalibrered on the source, after correction for the energy-encircled fraction. The background counts for AGNs are determined by summing all values from a radius circle centered at their respective coordinates in the modeled background map. Similarly, the exposure times for the AGNs are determined by reading the values from the exposure map at their respective coordinates. Figure 2 shows the coordinates of MIR-selected AGNs on the exposure map and their corresponding exposure times. Columns 5–6 of Table 2 give the aperture background counts and exposure time for each source.
Applying the calculation process described above, we derived X-ray luminosity upper limits for the entire MIR-selected AGN sample. Columns 7–8 of Table 2 list the upper limits of flux density and luminosity in the 2–10 keV band.